OpenStudy (anonymous):

Find the slope given the points: (-7,6), (-7,-9) Can u show me the work please!

OpenStudy (anonymous):

Where the slope of a line is m: \[m = (y_1-y_2)/(x_1-x_2)\] which is m = (6-9)/(-7-7) m = -3/-14 m = 3/14

OpenStudy (anonymous):

so helpful thanks :)